OpenStudy (anonymous):

what is the sine value of pi/6?

jimthompson5910 (jim_thompson5910):

Use the unit circle http://www.regentsprep.org/regents/math/algtrig/att5/600px-Unit_circle_angles_svg.jpg look for the angle pi/6 then look at the corresponding point on the unit circle. Specifically, look at the y coordinate since y = sin(theta)

OpenStudy (anonymous):

Ok!

OpenStudy (anonymous):

1/2?

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

if you wanted cos(pi/6), then that would be the x coordinate of that same point, so \[\Large \cos\left(\frac{\pi}{6}\right) = \frac{\sqrt{3}}{2}\]

jimthompson5910 (jim_thompson5910):

since x = cos(theta)

OpenStudy (anonymous):

Ohh! cAN I ask one more?

OpenStudy (anonymous):

Compare the following functions: f(x) = − 5 sin(2x − π) + 2 g(x) graph of a quadratic with points at 1, 2 and 3, negative 2 and 5, 2 h(x) x y −2 10 −1 7 0 5 1 3 2 5 3 7 4 10 Which function has the smallest minimum?

OpenStudy (anonymous):

@jim_thompson5910